For the following scores, find the mean, median, sum of squared deviations, variance and standard deviation. 2.8. 3.1. 2.6 3.3. 2.9 3.2
Answer #1

Mean =( 2.8+ 3.1+ 2.6+ 3.3.+ 2.9 + 3.2 )/6 = 2.98333

Median : First arrange the scores in incresing order

2.6 , 2.8 , 2.9 , 3.1 , 3.2 , 3.3

Now, Median = (2.9 + 3.1)/2 = 3

Sum of sqaured Deviation

= ( 2.8-2.98)2 + ( 3.1-2.98)2 + ( 2.6-2.98)2 + (3.3-2.98)2 + (2.9-2.98)2 +(3.2-2.98)2 = 0.3484

Variance = Sum of squared deviation/n

= 0.3484/6 = 0.058066

Standard deviation = √variance

   =√(0.058066) = 0.240970